OpenStudy (anonymous):

Brandon is 4 times as old as Kevin and is also 12 years older than Kevin.How old is Kevin?

OpenStudy (anonymous):

B=4*K Also B=K+12 so 4K=K+12 or K=4 years old

OpenStudy (anonymous):

thanks :)